Software Engineering

Testing stage of a project cost $977000 and took 8 months to complete. Estimate the overall project cost, the cost per year, its manpower, and the total number of LOC and the cost of each of the three remaining major effort allocations for the project. Assume the typical $k 100/year per average engineer. Check using COCOMO if these numbers are within reason.

Requirements is 24%

Design is 16%

Coding is 12%

Testing is 48%
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

Since testing is the well known 48% of the total, then

Project Total Cost = 977000/0.48 = $ 2035416.67

Project Total Time = 8/0.48 = 16.67 months.

the rest of the money and time is divided as per those % we learned in class!

Lifecycle Stage %  Cost [$] Time [months]

Requirements  24  488500.0008  4.0008

Design   16  325666.6672  2.6672

Coding 12 244250.0004  2.0004

Testing  48 977,000 8.00

TOTALS           100 2035416.67   16.67

But this cost is the total for 16.67 months.

Cost Per Year = 2035416.67*(12/16.67) = $ 1465206.961008

We divide this amount by 3.0 to look at real salaries: 1465206.961008/3 = $ 488402.320336

and this looks like just 488402.320336/100,000 = 4.884023 reasonably well paid Engineers!!!

4.884023 engineers at 1000 LOC/month produce = 4,884 LOC/month

[5 pts] Total LOCs in 2 (12% of 16.67) months: 4,884*2= 9768 LOC = 9.768 KLOC.

COCOMO verification:

The project has 4.884023 persons for 16.67 months =81.416663 (pm)

E = a(9.77)b = 3.6(9.77)1.20 = 55.485

Then 55.485/81.416663 = 68.14%. For a 70% estimate we need 55.485/.7 = 79.264,

79.264/16.67 = 4.75 engineers instead of 4.44 … QUITE GOOD! Our engineers are making just a bit more than the 100K$

Kindly revert for any queries

Thanks.